OpenStudy (anonymous):

I would really appreciate some help: When air expands adiabatically (without gaining or losing heat), its pressure P and volume V are related by the equation PV^1.4=C where C is a constant. Suppose that at a certain instant the volume is 340 cubic centimeters and the pressure is 83 kPa and is decreasing at a rate of 12 kPa/minute. At what rate in cubic centimeters per minute is the volume increasing at this instant?

OpenStudy (unklerhaukus):

\[P = 340 (cm^{3})=3.4\times10^{-4}(m^3)\] \[V=83(kPa) \] \[{dP \over dt} = 12 {(kPa/min)} = 0.2({kPa \over s})\]

OpenStudy (unklerhaukus):

\[V={nRT \over P}\]\[\ V^{1.4}=({nRT\over P})^{1.4} \] \[V^{1.4}=(nRT)^{1.4}P^{-1+1.4}\] \[V^{1.4} =C' P^{0.4 }\] --- im sorry but i cant figure out how to solve this problem (and im not sure if anything i have written is helpful)